OpenStudy (anonymous):

Which of the following is the solution to the equation 25^(z + 2) = 125 ? z = 5.5 z = 3.5 z = -2.5 z = -0.5

OpenStudy (callisto):

\[25^{z + 2} = 125\]\[5^{2(z + 2)} = 5^3\]\[2(z+2) =3\] Can you solve it?

OpenStudy (anonymous):

im not sure how? :/

OpenStudy (callisto):

multiply (z+2) by 2, what do you get of the left side?

OpenStudy (anonymous):

2z+4?

OpenStudy (callisto):

Yes, now you have 2x+4 =3 Can you solve this?

OpenStudy (anonymous):

i dont know the next step sorry

OpenStudy (callisto):

Subtract 4 from both sides, what do you get?

OpenStudy (anonymous):

2x=-1

OpenStudy (callisto):

Yes, now, divide both sides by 2, what do you get?

OpenStudy (anonymous):

-0.5

OpenStudy (callisto):

Yes :D

OpenStudy (anonymous):

thank you :D

OpenStudy (callisto):

Welcome!!!

OpenStudy (anonymous):

hey can i ask you something

OpenStudy (callisto):

Yes?

OpenStudy (anonymous):

i get all the steps but i was just wondering how u got 52(z+2)=53 2(z+2)=3 that part i didn't understand.

OpenStudy (callisto):

\[25 = 5 \times 5 = 5^2\]and \[125 = 5 \times 5 \times 5 =5^3\]

OpenStudy (anonymous):

ohhhhhhhh ok now i get it :]

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

And the question is \[25^{z+2} = 125\]So, \[(5^2)^{z+2} = 5^3\]By properties of exponential functions, \[5^{2(z+2)} = 5^3\]

OpenStudy (callisto):

Welcome!~